### v2

# Linear Resistivity from Spatially Random Interactions and the Uniqueness of Yukawa Coupling

Abstract
Recent studies have shown that a spatially random Yukawa-type interaction between a Fermi surface and critical bosons can produce linear-in-temperature resistivity, the defining signature of strange metals. In this article, we systematically classify all scalar couplings of the form $(\psi^{\dagger}\psi)^{n}\phi^{m}$ in arbitrary dimensions to identify possible candidates for strange-metal behaviour within this disordered framework. We find that only the Yukawa-type interaction in $(2+1)$ dimensions yields linear resistivity. Combined with previous results showing that spatially random QED-type couplings also produce such linearity, we conclude that spatially random Yukawa and QED-type interaction with external vector exhaust the class of random couplings capable of producing linear resistivity.
1 Introduction
Strange metals have emerged as one of the central topics in modern physics Lee2006 ; Anderson2017 ; Lee2018 ; Greene2020 ; Varma2020 ; Hartnoll2022 ; Phillips2022 . As the normal phase of superconductors, strange metals exhibit striking deviations from the Landau Fermi liquid theory. A defining feature of strange metals is linear-$T$ resistivity, $\rho\sim\rho_{0}+AT$, where $\rho_{0}$ is the residual resistivity and $A$ is a material-dependent coefficient. Despite extensive theoretical efforts over the past several decades, a systematic understanding of strange remains absent. A major obstacle is the lack of a theoretical framework accounting for the anomalous transport properties of strange metals, particularly the origin of linear resistivity.
More recently, Patel et al. Patel2023 introduced a spatially random coupling between a Fermi surface and a critical scalar boson, described by the interaction
$S_{\text{int}}=\int d\tau d^{2}{\bm{r}}\sum_{i,j,l=1}^{N}\frac{g_{ijl}({\bm{r}})}{N}\psi^{\dagger}_{i}(\tau,{\bm{r}})\psi_{j}(\tau,{\bm{r}})\phi_{l}(\tau,{\bm{r}}).$ (1)
The coupling constant $g_{ijl}({\bm{r}})$ follows a Gaußian distribution with zero mean and variance $\langle g_{ijl}^{*}({\bm{r}})g_{i^{\prime}j^{\prime}l^{\prime}}({\bm{r}}^{\prime})\rangle=g^{2}\delta({\bm{r}}-{\bm{r}}^{\prime})\delta_{ii^{\prime},jj^{\prime},ll^{\prime}}$. It was pointed out in Sin2025 that in this space dependent random coupling’s variance condition provides a wormhole picture in the field theory, which make it possible for far separated point to interact without distance dependent suppression. At large $N$ limit, this quenched disorder leads to linear resistivity at low temperatures. Subsequently, it was shown that the same strange metal transport property arises under an annealed average Sin2025 , and that the equivalence between quenched and annealed averages can be viewed as a concrete realisation of the $ER=EPR$ conjecture Maldacena2013 . This is different from the case where spatially homogeneous case which does not admit wormhole structure not the strange metalicity.
A similar fully randomised QED-like coupling between the critical Fermi surface and a vector boson can also reproduce such a linearity Wang:2024utm , even in the presence of a magnetic field Wang2025 . These all-to-all spatially random interactions, inspired by the Sachdev-Ye-Kitaev (SYK) model Sachdev1993 ; Kitaev2015 ; Chowdhury2022 , are referred to as ‘SYK-rised’ models in this article. Although such couplings cannot capture other anomalies of strange metals, such as Hall angles Wang2025 , they currently offer the only known mechanism that produces linear resistivity in a controlled, analytically tractable setting.
According to the analysis in ref.Wang:2024utm , Fermi’s Golden Rule can help to understand the emergence of linearity from the spatial random coupling (1), depite the absence of well-defined quasiparticles. In both scalar and vector models, the bosonic self-energy takes the form $\Pi(\Omega)\sim\Omega$, leading to a bosonic dispersion $\Omega\sim{\bm{q}}^{2}$. In $(d+1)$-dimensional systems, Bosonic density of states contributes a $T^{d/2}$-dependence to the resistivity. Meanwhile, the angular correction factor $(1-\cos\theta)$, with $\theta$ the scattering angle, is temperature-independent due to the relaxation of momentum conservation at each interaction vertex. Consequently, at low temperature, both interactions produce linear resistivity, though the linearity arises from different polarisation bubbles in Kubo formula Wang:2024utm .
SYK-rised models offer a promising route to understanding the nature of strange metals. This raises a natural question: can other types of random interactions between the FS and the critical boson also give rise to linear resistivity? In addition to higher-rank tensor couplings, one may consider interactions involving multiple fields. Specifically, we will investigate the SYK-rised interaction $(\psi^{\dagger}\psi)^{n}\phi^{m}$, and investigate the possibility that leads to linear-in-temperature resistivity across various spacetime dimensions.
This article is organised as follows. Section 2 develops a large-$N$ critical theory of spatial random couplings involving arbitrary numbers of fermions and bosons. The corresponding conductivity is calculated in section 3. We find that Yukawa-type couplings are the only class that produces linear-$T$ resistivity. Concluding remarks are given in Section 4.
2 Generalised Spatially Random Coupling
2.1 $G-\Sigma$ formalism
We begin by reviewing the model introduced in ref.Patel2023 , which provides the first theoretical realisation of linear-in-temperature resistivity. The fermionic field $\psi_{i}$ and scalar field $\phi_{i}$ are described by standard kintic terms,
$\mathcal{S}_{0}$ $=$ $\int d\tau\sum_{{\bm{k}}}\sum_{i=1}^{N}\psi^{\dagger}_{i{\bm{k}}}(\tau)\left[\partial_{\tau}+\varepsilon({\bm{k}})-\mu\right]\psi_{i{\bm{k}}}(\tau)+\int d\tau d^{2}{\bm{r}}\psi^{\dagger}_{i}(\tau,{\bm{r}})v_{ij}({\bm{r}})\psi_{j}(\tau,{\bm{r}})$ (2)
$+\frac{1}{2}\int d\tau\sum_{{\bm{q}}}\sum_{i=1}^{N}\phi_{i{\bm{q}}}(\tau)\left[-\partial_{\tau}^{2}+{\bm{q}}^{2}+m_{b}^{2}\right]\phi_{i,-{\bm{q}}}(\tau),$
where i=1,,Ni=1,...,N denotes the flacour index. The potential disorder $v_{ij}$ is introduced with Gausßian statistics,
$\langle v_{ij}({\bm{r}})\rangle=0\,,\quad\langle v^{\ast}_{ij}({\bm{r}})v_{lm}({\bm{r}}^{\prime})\rangle=v^{2}\,\delta({\bm{r}}-{\bm{r}}^{\prime})\delta_{il}\delta_{jm}.$ (3)
The key ingredient is a quenched disorder interaction between electrons and bosons,
$S_{\text{int}}=\int d\tau d^{2}{\bm{r}}\sum_{i,j,l=1}^{N}\frac{g_{ijl}({\bm{r}})}{N}\psi^{\dagger}_{i}({\bm{r}})\psi_{j}({\bm{r}})\phi_{l}({\bm{r}},\tau),$ (4)
with the coupling constants satisfying
$\langle g_{jil}({\bm{r}})\rangle=0,\quad\langle g^{*}_{ijl}({\bm{r}})g_{i^{\prime}j^{\prime}l^{\prime}}({\bm{r}}^{\prime})\rangle=g^{2}\delta_{ii^{\prime}jj^{\prime}ll^{\prime}}\delta({\bm{r}}-{\bm{r}}^{\prime}).$ (5)
We refer to the interaction (5) as an ‘SYK-rised’ Yukawa coupling, as it describes a spatially random all-to-all interaction analogous to the SYK model. It has been shown that in the large-$N$ limit, this interaction yields linear resistivity at low temperatures. Moreover, such SYK-rised mechanism remains effective when extended to vector bosons Wang:2024utm . This raises two natural questions: whether this is the unique scalar coupling that leads to linear resistivity, and whether such behaviour persists in higher-dimensional systems.
A natural generalisation of interaction (5) is to increase the number of fields involved. Specifically, we consider an interaction fo the form
$S_{g}$ $=$ $\frac{g_{\{i\}\{j\}\{l\}}({\bm{r}})}{\zeta}\int d\tau d^{d}r\sum_{\{i\},\{j\},\{l\}}\psi^{\dagger}_{i_{1},{\bm{k}}}({\bm{r}},\tau)...\psi^{\dagger}_{i_{n},{\bm{k}}}({\bm{r}},\tau)$ (6)
$\times$ $\psi_{j_{1},{\bm{k}}+{\bm{q}}}({\bm{r}},\tau)...\psi_{j_{n},{\bm{k}}+{\bm{q}}}({\bm{r}},\tau)\phi_{l_{1},{\bm{q}}}({\bm{r}},\tau)...\phi_{l_{m},{\bm{q}}}({\bm{r}},\tau),$
where $\zeta$ is a normalisation factor to be determined. The vertex involves $2n$ fermionic fields and $m$ bosonic fields, whilst the coupling constant $g_{\{i\}\{j\}\{l\}}\equiv g_{i_{1}...i_{n}j1...j_{n}l_{1}...l_{m}}$ obeys Gaußian distribution,
$\langle g_{\{i\}\{j\}\{l\}}({\bm{r}})\rangle=0,$ (7)
$\langle g_{\{i\}\{j\}\{l\}}g^{*}_{\{i^{\prime}\}\{j^{\prime}\}\{l^{\prime}\}}({\bm{r}}^{\prime})\rangle=g^{2}\delta({\bm{r}}-{\bm{r}}^{\prime})\delta_{\{i\}\{i^{\prime}\},\{j\}\{j^{\prime}\},\{l\}\{l^{\prime}\}},$ (8)
with multi-index delta functions defined by
$\delta_{\{i\}\{i^{\prime}\}}\equiv\delta_{i_{1}i_{1}^{\prime}}\delta_{i_{2}i_{2}^{\prime}}...\delta_{i_{n}i_{n}^{\prime}}.$ (9)
Instead of working in $(2+1)$ dimensions, we consider general (d+1)(d+1)D systems to find the effect of dimensionality. To determine the appropriate value of $\zeta$, we require both fermionic and bosonic self-energies remain of order $\mathcal{O}(1)$ in large-$N$ limit. This ensures that the theory is well controlled in IR Esterlis2021 ; Lee2009 . Under this condition, one obtains
$\zeta=N^{(2n+m-1)/2}.$ (10)
We now derive the $G-\Sigma$ action of theory (6). We define bilocal variables
$G(x_{1},x_{2})\equiv-\frac{1}{N}\sum_{i=1}^{N}\psi_{i}(x_{1})\psi_{i}^{\dagger}(x_{2}),$ (11)
$D(x_{1},x_{2})\equiv\frac{1}{N}\sum_{i=1}^{N}\phi_{i}(x_{1})\phi_{i}(x_{2}),$ (12)
and imposing them via Lagrange multipliers. This gives rise to the term
$S_{L}$ $=$ $-N\int d\tau d\tau^{\prime}\sum_{k}\Sigma({\bm{k}},\tau^{\prime}-\tau)\left[G({\bm{k}},\tau-\tau^{\prime})+\frac{1}{N}\sum_{i=1}^{N}\psi_{i,{\bm{k}}}(\tau)\psi^{\dagger}_{i,{\bm{k}}}(\tau^{\prime})\right]$ (13)
$+\frac{N}{2}\int d\tau d\tau^{\prime}\sum_{q}\Pi({\bm{q}},\tau^{\prime}-\tau)\left[D({\bm{q}},\tau-\tau^{\prime})-\frac{1}{N}\sum_{i=1}^{N}\phi_{i,{\bm{q}}}(\tau)\phi_{i,-{\bm{q}}}(\tau^{\prime})\right].$
Following the analysis in Ref.Esterlis2021 , we neglect the replica off-diagonal components of $G(x_{1},x_{2})$ and $D(x_{1},x_{2})$ in large $N$. As emphasised in ref.Cardy1996 ; Esterlis2021 , the universal properties of critical theories are expected to be insensitive to the microscopic details of the interaction. We therefore assume replica symmetry from the outset.
Then we perform the disorder average using the replica trick. The partition function is given by
$\mathcal{Z}\equiv\int\mathcal{D}[\psi,\psi^{\dagger}]\mathcal{D}[\phi]e^{-S_{0}-S_{g}-S_{\text{L}}},$ (14)
and the resulting Gaußian integrals can be evaluated as
$\int\mathcal{D}[\psi,\psi^{\dagger}]e^{-\psi^{\dagger}\mathbf{A}\psi}=\det\mathbf{A},$ (15)
$\int\mathcal{D}[\phi]e^{-\phi\mathbf{A}\phi/2}=\det\mathbf{A}^{-1/2}.$ (16)
Applying these to the replicated action yields the effective action
$\frac{S[G,\Sigma;D,\Pi]}{N}$ $=$ $-\ln\det(\partial_{\tau}+\varepsilon({\bm{k}})\delta(x-x^{\prime})-\mu+\Sigma)$ (17)
$+\frac{1}{2}\ln\det\left(\left(-\partial_{\tau}^{2}+K{\bm{q}}^{2}+m_{b}^{2}\right)\delta(x-x^{\prime})-\Pi\right)$
[…]
where $\bar{\delta}$ is a Dirac delta over space, $\delta({\bm{r}}-{\bm{r}}^{\prime})$, and the trace notation is defined by Gu2020
$Tr(f\cdot g)\equiv f^{T}g\equiv\int dx_{1}dx_{2}f(x_{2},x_{1})g(x_{1},x_{2})\,.$ (18)
Assuming replica symmetry, we drop replica indices without loss of generality.
2.2 Large-$N$ Critical Theory
The large-$N$ limit is governed by the saddle point of the $G-\Sigma$ action (17). Varying the action yields
$0=\frac{\delta S}{N}$ (19)
[…]
Using the identity
[…]
we obtain the saddle-point equations (Schwinger-Dyson equations),
$G_{*}[\Sigma](x_{1},x_{2})$ $=$ $(-\partial_{\tau}-\varepsilon({\bm{k}})+\mu-\Sigma)^{-1}(x_{1},x_{2})\,$, (21)
$\Sigma_{*}[G](x_{1},x_{2})$ $=$ $ng^{2}G^{2n-1}(x_{1},x_{2})D^{m}(x_{1},x_{2})\bar{\delta}+v^{2}G(x_{1},x_{2})\bar{\delta},$ (22)
$D_{*}[\Pi](x_{1},x_{2})$ $=$ $(-\partial_{\tau}^{2}+K{\bm{q}}^{2}+m_{b}^{2}-\Pi)^{-1}(x_{1},x_{2}),$ (23)
$\Pi_{*}[D](x_{1},x_{2})$ $=$ $mg^{2}G^{2n}(x_{1},x_{2})D^{m-1}(x_{1},x_{2})\bar{\delta}.$ (24)
Setting $n=m=1$ recovers the saddle-point structure of the SYK-rised Yukawa model discussed in Refs.Patel2023 ; Esterlis2021 ; Guo2022 . We adopt the conventional quadratic dispersion $\varepsilon({\bm{k}})={\bm{k}}^{2}/(2m)$, and fix units such that the boson velocity $\sqrt{K}=1$.
The next step is to solve the Schwinger-Dyson equations obtained above at $T=0$. To simplify the calculation, we define $\xi_{{\bm{k}}}\equiv\varepsilon({\bm{k}})-\mu$. The contribution from potential disorder $v_{ij}({\bm{r}})$ is straightforward to evaluate,
$\Sigma_{v}(\Omega_{m})$ $=$ $v^{2}\int\frac{d^{d}{\bm{k}}}{(2\pi)^{2}}G(i\Omega_{m},{\bm{k}})$ (25)
$=$ $v^{2}{\mathcal{N}}\int d\xi_{{\bm{k}}}\frac{1}{i\Omega_{m}-\xi_{{\bm{k}}}-\Sigma(i\Omega_{m},{\bm{k}})}$
$=$ $-i\frac{\Gamma}{2}{\mathrm{sgn}}(\Omega_{m}),$
[…]
We now assume that impurity scattering dominates over interactions, i.e. $|\Sigma_{v}|\gg|\Sigma_{g}|$, so that the electron propagator is well approximated by potential disorder alone. At low frequencies,
$G({\mathrm{i}}\omega,{\bm{k}})$ $\simeq$ $\frac{1}{{\mathrm{i}}{\mathrm{sgn}}(\omega)\Gamma/2-{\bm{k}}^{2}/(2m)+\mu},$ (26)
where the corrections from fermion-boson coupling are suppressed at leading order.
At critical point, the boson mass satisfies $m_{b}^{2}-\Pi(0,0)=0$ Esterlis2021 . The full bosonic propagator can be expressed as
$D({\mathrm{i}}\Omega_{m},{\bm{q}})$ $=$ $\frac{1}{\Omega_{m}^{2}+{\bm{q}}^{2}+m_{b}^{2}-\Pi(\Omega_{m}=0)-(\Pi-\Pi(\Omega_{m}=0))}$ (27)
$=$ $\frac{1}{\Omega_{m}^{2}+{\bm{q}}^{2}-(\Pi-\Pi(\Omega_{m}=0))}.$
It remains to evaluate $\Pi({\mathrm{i}}\Omega)-\Pi(0,0)$, which encodes the leading dynamical correction to the critical bosonic propagator.
We now turn to the evaluation of the self-energies. The boson self-energy (24) is graphically represented by Fig.1(a), where solid lines denote fermion propagators, wavy lines denote bosons, and dashed lines indicate disorder averaging. The fermionic self-energy (22) is illustrated in Fig.1(b). Since bosonic self-energy and bosonic self-energy are structurally similar, we do not need to compute them separately. Instead, we can analyse both simultaneously by studying the frequency dependence of a generic diagram containing $\alpha$ fermion lines and $\beta$ boson lines.  [^1] A non-zero $\beta$ requires $m\geq 2$, but the scaling behaviour obtained below remains valid even for $m=1$; further details are provided in present in Appendix A.
(a) Bosonic self-energy
(b) Fermionic self-energy
Figure 1: Two self-energies, $\Pi$ and $\Sigma$, share the same structure.
Two integrals will be useful in evaluating the generic diagram. First, we note that
$\int\frac{x^{d-1}}{A+x^{2}}dx=\frac{x^{d}\,_{2}F_{1}\left(1,\frac{d}{2};\frac{d}{2}+1;-\frac{x^{2}}{A}\right)}{Ad},$ (28)
where $_{2}F_{1}(a,b;c;d)$ is the hypergeometric function. This leads to the estimate
$\int_{0}^{\infty}\frac{x^{d-1}}{A+x^{2}}dx\sim A^{\frac{d-2}{2}},$ (29)
where we retain only the dependence on $A$, which is sufficient for the scaling analysis below. The second integral is
$\int\left(\prod_{i=1}^{n}dx_{i}{\mathrm{sgn}}(x_{i})\right)\Big{|}B-\sum_{i}^{n}x_{i}\Big{|}^{\lambda}$ (30)
$\sim$ $\underset{|\sum_{i}^{n}x_{i}|\leq|B|}{\int_{0}^{\infty}}\prod_{i=1}^{n}dx_{i}\Big{|}B-\sum_{i}^{n}x_{i}\Big{|}^{\lambda}=\underset{|\sum_{i}^{n}x_{i}|\leq|B|}{\int_{0}^{\infty}}\prod_{i=1}^{n}dx_{i}\Big{|}B-\sum_{i}^{n}x_{i}\Big{|}^{\lambda}\int dX\delta(X-\sum_{i}^{n}x_{i})$
$\sim$ $\int_{0}^{|B|}d|X|\prod_{i=1}^{n-1}\int_{0}^{|X|}dx_{i}\Big{|}B-X\Big{|}^{\lambda}=\int_{0}^{B}dX|X|^{n-1}\Big{|}B-X\Big{|}^{\lambda}$
$\sim$ $|B|^{n+\lambda},$
where we have defined $X\equiv\sum_{i}^{n}x_{i}$, and inserted a delta function to isolate the constraint.
The fermionic Green’s function is approximately given by eqn.(26), but bosonic propagator remains undetermined. As a result, generic diagrams with $\beta\geq 1$ appear intractable. However, they can still be analysed by assuming a scaling form for the bosonic self-energy,
$\Pi({\mathrm{i}}\Omega_{m})-\Pi(0)\sim-c_{B}|\Omega_{m}|^{\eta^{\prime}},$ (31)
where $\eta^{\prime}$ will be fixed self-consistently at the end of the calculation. At low frequencies, if $\eta<2$, the bosonic propagator reads
$D({\mathrm{i}}\Omega_{m},{\bm{q}})\simeq\frac{1}{{\bm{q}}^{2}+c_{B}|\Omega_{m}|^{\eta^{\prime}}}.$ (32)
Otherwise, the frequency dependence from self-energies is subleading, and
$D({\mathrm{i}}\Omega_{m},{\bm{q}})\simeq\frac{1}{\Omega_{m}^{2}+{\bm{q}}^{2}}.$ (33)
In either case, the propagator may be written uniformly as
$D({\mathrm{i}}\Omega_{m},{\bm{q}})\simeq\frac{1}{{\bm{q}}^{2}+c_{B}|\Omega_{m}|^{\eta}},$ (34)
with the requirement that $\eta\leq 2$.
Combining Eqs. (29) and (30), we evaluate the scaling behaviour of the generic diagram,
$\mathcal{I}^{d}_{\alpha,\beta}({\mathrm{i}}x)$ $\equiv$ $\int_{-\infty}^{\infty}\left(\prod_{i=1}^{\alpha}\frac{d\omega_{i}}{2\pi}d\xi_{{\bm{k}}_{i}}\frac{1}{{\mathrm{i}}{\mathrm{sgn}}(\omega_{i})\Gamma/2-\xi_{{\bm{k}}_{i}}}\right)$ (35)
$\times\left(\int_{-\infty}^{\infty}\prod_{j=1}^{\beta-1}\frac{d\Omega_{j}}{2\pi}\int_{0}^{\infty}\frac{d^{d}{\bm{q}}_{j}}{(2\pi)^{2}}\frac{1}{c_{B}|\Omega_{j}|^{\eta}+{\bm{q}}_{j}^{2}}\right)\int_{0}^{\infty}\frac{d^{d}{\bm{q}}_{\beta}}{(2\pi)^{2}}\frac{1}{c_{B}|x+\sum_{j=1}^{\beta-1}\Omega_{j}+\sum_{i=1}^{2n}\omega_{i}|^{\eta}+{\bm{q}}_{\beta}^{2}}$
$\sim$ $(-{\mathrm{i}})^{\alpha}\int_{-\infty}^{\infty}\left(\prod_{i=1}^{\alpha}d\omega_{i}{\mathrm{sgn}}(\omega_{i})\right)\left(\prod_{j=1}^{\beta-1}\int_{-\infty}^{\infty}d\Omega_{j}|\Omega_{j}|^{\frac{\eta(d-2)}{2}}\right)\left|x+\sum_{j=1}^{\beta-1}\Omega_{j}+\sum_{i=1}^{2n}\omega_{i}\right|^{\frac{\eta(d-2)}{2}}$
$\sim$ $(-{\mathrm{i}})^{\alpha}\int_{-\infty}^{\infty}\left(\prod_{j=1}^{\beta-1}d\Omega_{j}|\Omega_{j}|^{\frac{\eta(d-2)}{2}}\right)\left|x+\sum_{j=1}^{\beta-1}\Omega_{j}\right|^{\frac{\eta(d-2)}{2}+\alpha}.$
We now perform the change of variables $\Omega_{j}\equiv x\cdot u_{j}$ so that $d\Omega_{j}=xdu_{j}$. The integral (35) then scales as
$\mathcal{I}^{d}_{\alpha,\beta}({\mathrm{i}}x)$ $\sim$ $(-{\mathrm{i}})^{\alpha}x^{\alpha+\beta+\frac{\eta(d-2)}{2}-1+\frac{\eta(d-2)(\beta-1)}{2}}.$ (36)
For bosonic self-energy, we identify $\alpha=2n$ and $\beta=m-1$; for fermionic self-energy, $\alpha=2n-1$ and $\beta=m$. This yields
$\Pi({\mathrm{i}}\Omega)-\Pi(0)$ $\sim$ $-c_{B}g^{2}\Omega^{2n+m-2+\frac{\eta(d-2)(m-1)}{2}},$ (37)
$\Sigma({\mathrm{i}}\omega)$ $\sim$ $-{\mathrm{i}}c_{F}g^{2}\omega^{2n+m-2+\frac{\eta(d-2)m}{2}},$ (38)
where $c_{B}$ and $c_{F}$ constants not relevant for scaling analysis.
Now let us determine $\eta$ by imposing self-consistency on the assumed scaling, $\Pi({\mathrm{i}}\Omega_{m})-\Pi(0)\sim-c_{B}|\Omega_{m}|^{\eta^{\prime}}$.
  • case i: If $\eta^{\prime}<2$, then $\eta=\eta^{\prime}$, and $\eta$ satisfies
    $2n+m-2+\frac{\eta(d-2)}{2}+\frac{\eta(d-2)(m-2)}{2}=\eta.$ (39)
    Solving for $\eta$, one finds
    $\eta=-\frac{2(m+2n-2)}{dm-d-2m}.$ (40)
    Substituting this back into the self-energies (37) and (38), we have
    $\Pi({\mathrm{i}}\Omega)-\Pi(0)$ $\sim$ $-c_{B}g^{2}\Omega^{-\frac{2(m+2n-2)}{d(m-1)-2m}},$ (41)
    $\Sigma({\mathrm{i}}\omega)$ $\sim$ $-{\mathrm{i}}c_{F}g^{2}\omega^{-\frac{d(m+2n-2)}{d(m-1)-2m}}.$ (42)
    For this solution to be consistent, the exponent must satisfy
    $-\frac{2(m+2n-2)}{d(m-1)-2m}<2.$ (43)
  • case ii: If $\eta^{\prime}\geq 2$, then $\eta=2$, yielding
    $\Pi({\mathrm{i}}\Omega)-\Pi(0)$ $\sim$ $-c_{B}g^{2}\Omega^{2n+m-2+(d-2)(m-1)},$ (44)
    $\Sigma({\mathrm{i}}\omega)$ $\sim$ $-{\mathrm{i}}c_{F}g^{2}\omega^{2n+m-2+(d-2)m},$ (45)
    with the consistency condition
    $2n+m-2+(d-2)(m-1)\geq 2.$ (46)
3 Conductivity
3.1 The unique approach to linearity
Having obtained the propagators and self-energies, we now estimate the conductivity using the Kubo formula,
$\sigma^{\mu\nu}({\mathrm{i}}\Omega_{m})$ $=$ $-\frac{1}{\Omega_{m}}\left[\langle j^{\mu}(\Omega)j^{\nu}(-\Omega)\rangle\right]_{\Omega=0}^{\Omega={\mathrm{i}}\Omega_{m}}$ (47)
$\equiv$ $-\frac{1}{\Omega_{m}}\left[\tilde{\Pi}^{\mu\nu}(\Omega)\right]_{\Omega=0}^{\Omega={\mathrm{i}}\Omega_{m}}$
where $\tilde{\Pi}^{\mu\nu}$ denotes the current-current correlator, or equivalently, the polarisation bubble for the external electromagnetic field. Since we consider only electric fields in this work, the polarisation tensor, so $\tilde{\Pi}^{\mu\nu}$ is diagonal.
(a) The simplest polarisation bubble.
(b) Self-energy correction.
(c) MT diagram
(d) AL diagrams
Figure 2: Polarisation diagrams. The solid line represents the “bare” propagator (26).
The polarisation bubble receives contributions from the simplest one-loop diagram, as well as from self-energy insertions and vertex corrections, as illustrated in Fig. 2. To leading order, the polarisation is illustrated in Fig. 2(a), and is given by Coleman2019
$\tilde{\Pi}^{\mu\nu}_{0}({\mathrm{i}}\Omega_{m})$ (48)
$=$ $-\frac{2}{m^{2}}T\sum_{n}\int\frac{d^{d}{\bm{k}}}{(2\pi)^{2}}k^{\mu}k^{\nu}\frac{1}{{\mathrm{i}}\omega_{n}-\frac{{\bm{k}}^{2}}{2m}+\mu+{\mathrm{i}}\Gamma{\mathrm{sgn}}(\omega_{n})/2}\frac{1}{{\mathrm{i}}(\omega_{n}+\Omega_{m})-\frac{{\bm{k}}^{2}}{2m}+\mu+{\mathrm{i}}\Gamma{\mathrm{sgn}}(\omega_{n}+\Omega_{m})/2}$
$\simeq$ $-\frac{k_{F}^{2}\mathcal{N}}{m^{2}}\delta^{\mu\nu}\int\frac{d\omega_{n}}{2\pi}d\xi_{{\bm{k}}}\frac{1}{{\mathrm{i}}\omega_{n}-\xi_{{\bm{k}}}+{\mathrm{i}}\Gamma{\mathrm{sgn}}(\omega_{n})/2}\frac{1}{{\mathrm{i}}(\omega_{n}+\Omega_{m})-\xi_{{\bm{k}}}+{\mathrm{i}}\Gamma{\mathrm{sgn}}(\omega_{n}+\Omega_{m})/2}$
$=$ $-v_{F}^{2}\mathcal{N}\delta^{\mu\nu}\frac{\Omega_{m}}{\Omega_{m}+{\mathrm{sgn}}(\Omega_{m})\Gamma},$
which leads to a finite constant residual resistivity. Here we have assumed that the dominant contributions come from fermions near the Fermi surface Coleman2019 .
At the next order, the polarisation receives corrections from the fermionic self-energy, shown in Fig. 2(b). For convenience, we write the self-energy in the form
$\Sigma({\mathrm{i}}\omega)\equiv-{\mathrm{i}}c_{F}g^{2}\omega^{\varsigma}.$ (49)
The corresponding correction to the polarisation tensor is
$\tilde{\Pi}_{g}^{\mu\nu}({\mathrm{i}}\Omega)$ $=$ $-\frac{2}{m^{2}}\int_{-\infty}^{\infty}\frac{d\omega}{2\pi}\int_{-\infty}^{\infty}\frac{d^{2}{\bm{k}}}{(2\pi)^{2}}k^{\mu}k^{\nu}G({\mathrm{i}}\omega_{n})\Sigma({\mathrm{i}}\omega)G({\mathrm{i}}\omega)G({\mathrm{i}}(\omega+\Omega))$ (50)
$\sim$ $-\frac{2k_{F}^{2}\mathcal{N}}{m2}\delta^{\mu\nu}\int\frac{d\omega_{n}}{2\pi}d\xi_{{\bm{k}}}\frac{1}{{\mathrm{i}}\omega_{n}-\xi_{{\bm{k}}}+{\mathrm{i}}\Gamma{\mathrm{sgn}}(\omega)/2}\frac{1}{{\mathrm{i}}\omega_{n}-\xi_{{\bm{k}}}+{\mathrm{i}}\Gamma{\mathrm{sgn}}(\omega)/2}$
$\times\frac{1}{{\mathrm{i}}(\omega+\Omega)-\xi_{{\bm{k}}}+{\mathrm{i}}\Gamma{\mathrm{sgn}}(\omega+\Omega)/2}\left(-{\mathrm{i}}c_{F}g^{2}\omega^{\varsigma}\right)$
$\sim$ $-g^{2}\delta^{\mu\nu}\int d\omega\frac{{\mathrm{sgn}}(\omega)-{\mathrm{sgn}}(\omega+\Omega)}{\left({\mathrm{i}}\frac{\Gamma}{2}{\mathrm{sgn}}(\omega)-{\mathrm{i}}\frac{\Gamma}{2}{\mathrm{sgn}}(\omega+\Omega)\right)^{2}}\omega^{\varsigma}$
$\sim$ $g^{2}\delta^{\mu\nu}\Omega^{\varsigma+1}$
at low frequencies.
The vertex corrections—specifically, the Maki-Thompson (MT) diagram (Fig.2(c)) and the Aslamazov-Larkin (AL) diagrams (Fig.2(d)), vanish in our model. This cancellation arises because the disorder average enforces spatial locality, effectively decoupling all momenta in the internal propagators via a delta function in position space Patel2023 . Take the MT diagram as an example: after disorder averaging, it yields
$\tilde{\Pi}_{\text{MT}}({\mathrm{i}}\Omega)\sim\int d^{2}{\bm{k}}d^{2}{\bm{k}}^{\prime}{\bm{k}}{\bm{k}}^{\prime}G({\bm{k}},{\mathrm{i}}\omega)G({\bm{k}},{\mathrm{i}}(\omega+\Omega))G({\bm{k}}^{\prime},{\mathrm{i}}\omega^{\prime})G({\bm{k}}^{\prime},{\mathrm{i}}(\omega^{\prime}+\Omega))$..., (51)
where the integrand is odd under both ${\bm{k}}\to-{\bm{k}}$ and ${\bm{k}}^{\prime}\to-{\bm{k}}^{\prime}$, and thus integrates to zero. The same argument applies to the AL diagrams. As a result, all vertex corrections vanish in these SYK-rised models.
Having computed all relevant polarisation tensors, we now apply the Kubo formula to extract the resistivity. Denoting the self-energy correction to the polarisation as $\tilde{\Pi}_{g}\equiv-c_{g}g^{2}\Omega^{1+\varsigma}$, where $c_{g}$ is a constant, the total polarisation tensor to order $\mathcal{O}(g^{2})$ reads
$\tilde{\Pi}^{\mu\nu}\simeq\tilde{\Pi}^{\mu\nu}_{0}+\tilde{\Pi}^{\mu\nu}_{g}.$ (52)
Substituting into the Kubo formula (47) and analytically continuing (${\mathrm{i}}\Omega\to\Omega$), we obtain the conductivity
$\sigma^{\mu\nu}(\Omega)=v_{F}^{2}\mathcal{N}\delta^{\mu\nu}\frac{1}{\Gamma-{\mathrm{i}}\Omega}-c_{g}g^{2}\delta^{\mu\nu}(-{\mathrm{i}}\Omega)^{\varsigma}.$ (53)
In the absence of a magnetic field, the tensor structure is trivial and $\tilde{\Pi}$ is diagonal, so we drop the indices and write the scalar conductivity. Taking the inverse and extracting the real part, the resistivity is
[…]
where $c_{g}^{\prime}$ absorbs numerical prefactors independent of $g^{2}$.
The frequency dependence at zero temperature translates to $T$-dependence at finite temperature. Our goal is to identify all combinations of $d,m,n$ that yield a linear resistivity. For $\eta<2$, it corresponds to solving
$d(m+2n-2)=2m-d(m-1),$ (55)
which can be rearranged as
$n=\left(\frac{1}{d}-1\right)m+\frac{3}{2}\in\mathbb{Z}^{+}.$ (56)
There is no solution for $d=1$, so we focus on $d\geq 2$. As $n,m\in\mathbb{Z}^{+}$, we also impose
$0<m\leq\frac{1}{2}\frac{d}{d-1}.$ (57)
The only integer $m$ satisfying this is $m=1$ when $d=2$. Substituting $d=2$ and $m=1$ into eqn.(56) gives $n=1$. One can further verify that when $n=m=1$ and $d=2$, $\Pi\sim\Omega$, consistent with the assumption.
For $\eta=2$, we need to solve
$\varsigma=2n+m-2+(d-2)m=1.$ (58)
No integer $n$ can satisfy this equation if $d=1$. When $d=2$, the only solution is $n=m=1$, but this leads to $\Pi\sim\Omega$, implying $\eta^{\prime}=1$, which contradicts the assumption $\eta^{\prime}\geq 2$. Thus, it cannot be a valid solution for $\eta=2$. For higher dimensions, since $2n-2\geq 0$ and $m+(d-2)m\geq 2m\geq 2$, he left-hand side of eqn.(58) is always greater than or equal to $2$. Consequently, there are no consistent integer solutions yielding linear resistivity for $\eta=2$.
Readers may notice that eqns.(53) and (54) differ from the results in ref.Patel2023 , where a logarithmic term $\ln(\Omega)$ appears but is absent here. This apparent discrepancy originates from the scaling approximation used in integral (29), which cannot capture logarithmic corrections. Instead, it yields a constant term when $d=2$. A direct, detailed calculation, rather than a general scaling analysis, naturally recovers the $\ln(\Omega)$ behaviour. Such a calculation for the case $d=2$ is presented in Appendix B.
In summary, the only consistent choice yielding linear resistivity is $d=2$ with $n=m=1$, precisely corresponding to the Yukawa-type coupling studied in ref. Patel2023 .
3.2 A detour: the failure of Fermi’s golden rule
Let us detour to the failure of Fermi’s golden rule in SYK-rised models. The bosonic dispersion relation takes the form
$\Omega\sim{\bm{q}}^{2/\eta}\equiv{\bm{q}}^{\alpha}.$ (59)
According to Fermi’s golden rule, the lifetime of an electron is estimated as
$\frac{1}{\tau}$ $\sim$ $\int\prod_{i=1}^{2n-1}d^{d}{\bm{k}}_{i}\,f_{{\bm{k}}_{i}}\prod_{j=1}^{m}d^{d}{\bm{q}}_{j}\,b_{{\bm{q}}_{j}}(1-\cos\theta)\,\delta\left(\sum_{i=1}^{2n-1}\omega_{i}+\sum_{j=1}^{m}\Omega_{j}-\omega_{2n}\right)$
$\sim$ $\int\prod_{i=1}^{2n-1}d\omega_{i}\,\frac{1}{e^{\beta\omega_{i}}+1}\prod_{j=1}^{m}d\Omega_{j}\,\frac{1}{e^{\beta\Omega_{j}}-1}\Omega_{j}^{\frac{d}{\alpha}-1}(1-\cos\theta)\,\delta\left(\sum_{i=1}^{2n-1}\omega_{i}+\sum_{j=1}^{m}\Omega_{j}-\omega_{2n}\right),$
where $f_{{\bm{k}}}$ and $b_{{\bm{k}}}$ are the Fermi-Dirac and Bose-Einstein distribution functions. In second step, we have changed variables from momenta to frequencies using despersion relation (59). Rescaling variables as $x_{i}\equiv\beta\omega_{i}$, $y_{j}\equiv\beta\Omega_{j}$, and $z=\beta\omega_{2n-1}$, we find
$\frac{1}{\tau}_{\text{GR}}$ $\sim$ $T^{m\frac{d}{\alpha}+2n-2}\int\prod_{i=1}^{2n-2}dx_{i}\,\frac{1}{e^{x_{i}}+1}\prod_{j=1}^{m}dy_{j}\,\frac{1}{e^{y_{j}}-1}\frac{1}{e^{z}+1}(1-\cos\theta).$ (61)
Meanwhile, the resistivity (54) derived from the Kubo formula scales as
$\rho_{\text{Kubo}}\sim T^{\varsigma}$ (62)
(from eqns. (42) and (45)), where
$\varsigma=\left\{\begin{aligned} -\frac{d(m+2n-2)}{d(m-1)-2m}=\frac{d}{\alpha},&\quad\eta^{\prime}<2\\ 2n+m-2+(d-2)m,&\quad\eta^{\prime}\geq 2\end{aligned}\right$. (63)
and
$\alpha=\left\{\begin{aligned} -\frac{d(m-1)-2m}{m+2n-2},&\quad\eta^{\prime}<2\\ 1,&\quad\eta^{\prime}\geq 2\end{aligned}\right$. (64)
as implied by the bosonic dispersion from (41). We see that the golden rule fails to reproduce the correct scaling: the result (61) does not match the Kubo-derived scaling (63), except in the special case $m=n=1$, where the two estimates coincide Wang:2024utm . Remarkably, this agreement occurs despite the absence of well-defined quasiparticles in two dimensions.
Admittedly, existence of well-defined quasiparticles, characterised by $\Sigma\sim\omega^{\varsigma}$ with $\varsigma\geq 2$, is a necessary condition for the applicability of the golden rule. However, the golden rule fails in SYK-rised models even when quasiparticles are present. In fact, there are no well-defined quasiparticles when $\eta<2$. To see this, consider the case with $n+m>2$, and distinguish two regimes within $\eta<2$:
  • i.
    If $\eta^{\prime}=\eta<0$, then $\varsigma=\eta d/2<0<2$. This outlaws the existence of quasiparticles.
  • ii.
    For $0<\eta^{\prime}=-2(m+2n-2)/[d(m-1)-2m]<2$, the condition
    $0<d<\frac{m-2n+2}{m-1}\leq\frac{m}{m-1}\leq 2,$ (65)
    is required, with $m>1$. If $m=1$, $\eta<2$ can only be satisfied by $n=1$, which violates the assumption $n+m>2$. Since $d<2$ in this regime, it follows that
    $\varsigma=\frac{\eta d}{2}<\eta=2,$ (66)
    indicating no quasiparticles.
again ruling out well-defined quasiparticles. Thus, in the regime $\eta<2$, the absence of quasiparticles can naturally explains why the golden rule fails.
On the other hand, when $\eta^{\prime}>2$, quasiparticles may exist. In this case, the fermionic exponent is $\varsigma=\eta^{\prime}+d-2$, since $\eta^{\prime}=2n+m-2+(d-2)(m-1)$ according to (44). If $\varsigma\geq 2$ (e.g.m=n=2m=n=2and $d=3$), the quasiparticles are well-defined. However, even in this regime, the golden rule does not yield the correct scaling of resistivity.
As discussed in the previous section, momentum conservation is relaxed at each interaction vertex due to disorder averaging. This decouples spatial momentum across different interaction lines. It is therefore highly likely that the golden rule, which relies on well-defined scattering kinematics, fails to correctly estimate the resistivity in SYK-rised models, even in the presence of quasiparticles.
4 Conclusion
In this work, we have systematically explored the full class of SYK-rised interactions of the form $(\psi^{\dagger}\psi)^{n}\phi^{m}$, involving multiple fermions and scalar bosons, as potential candidate theories for strange metals, across arbitrary dimensions. It turns out that only Yukawa-type interactions ($n=m=1$) in $2$ dimensions lead to linear-$T$ resistivity. There is no other candidate of strange metals within this SYK-rised class. Our analysis reveals that only the Yukawa-type coupling with $n=m=1$ in two spatial dimensions yields a linear-in-temperature resistivity. No other combination within this interaction class leads to such scaling, making the 2D Yukawa model the unique candidate among scalar-coupled SYK-rised theories.
In earlier work Wang:2024utm , we also demonstrated that a spatially random QED-like vector coupling, $\psi^{\dagger}\nabla^{\mu}\psi{\bm{a}}_{\mu}$, yields linear resistivity. Here $\bm{a}$} is a bosonic vector field, not necessarily a gauge field. Taken together, these results suggest that in realistic two-dimensional systems, the random Yukawa and vector couplings serve as minimal building blocks for capturing linear-$T$ resistivity. For any interaction of the form (6), the low-temperature resistivity scales as $\rho\sim T^{m+2n-2}$; thus, the linear case arises only when $n=m=1$, or equivalently from the vector coupling scenario. All higher order terms do not contribute to the linear resistivity. Therefore, there are essentially one class in each category: unique scalar class and unique vector class.
Beyond resistivity scaling, our study also uncovers the breakdown of Fermi’s Golden Rule in disordered models, even in cases where quasiparticles remain well-defined. Intriguingly, the only situation where the golden rule yields the correct resistivity scaling is again the Yukawa case ($n=m=1$), where quasiparticles are in fact absent. The mechanisms governing when the golden rule applies or fails in SYK-rised models remain unknown and deserve further investigation.
This paper has focused exclusively on low-temperature linear resistivity in SYK-rised models. While the Yukawa coupling emerges as a viable theory of strange metals within this regime, a more complete understanding requires going beyond the infrared limit. Notably, strange metals also exhibit linear resistivity at high temperatures, a regime not captured by our current theory. It remains an open question whether SYK-rised models can also account for $H$-linear resistivity and magnetoresistance obeying $H/T$ scaling Jiang2023 . These directions will be crucial for determining whether SYK-rised interactions offer a unified framework for understanding the strange metal phase.
Acknowledgements.
The authors would like to thank Ki-Seok Kim, Kyoung-Min Kim, Chao-Jung Lee, Sung-Sik Lee, Yi Zhang for the inspiring discussion. This work is supported by the National Research Foundation of Korea (NRF) grant funded by the Korean government (MSIT) No.NRF-2021R1A2B5B02002603, and No.RS-2023-00218998.
Appendix A Self-energies for $m=1$
As noted in the main text, the integral expression in eqn.(35), and hence the resulting self-energies in eqns.(37) and (38), are formally derived under the assumption $m\geq 2$. TIn this appendix, we explicitly compute the case $m=1$ to verify that these expressions remain valid even for $m=1$.
The bosonic self-energy in $(d+1)$ dimensions takes the form
$\Pi({\mathrm{i}}\Omega)-\Pi(0)$ (67)
$\simeq$ $-g^{2}m\mathcal{N}^{2n}\int\prod_{i=1}^{2n-1}\frac{d\omega_{i}}{2\pi}d\xi_{{\bm{k}}_{i}}d\xi_{{\bm{k}}_{2n}}\frac{1}{{\mathrm{i}}{\mathrm{sgn}}(\omega_{i})\Gamma/2-\xi_{{\bm{k}}_{i}}}$
$\times\left[\frac{1}{{\mathrm{i}}{\mathrm{sgn}}(\Omega+\sum_{i=1}^{2n-1}\omega_{i})\Gamma/2-\xi_{{\bm{k}}_{2n}}}-\frac{1}{{\mathrm{i}}{\mathrm{sgn}}(\sum_{i=1}^{2n-1}\omega_{i})\Gamma/2-\xi_{{\bm{k}}_{2n}}}\right]$
$\sim$ $-\frac{g^{2}m\mathcal{N}^{2n}}{2^{2n}}2\pi\int\prod_{i=1}^{2n-1}d\omega_{i}{\mathrm{sgn}}(\omega_{i})\left({\mathrm{sgn}}\left(\Omega+\sum_{i=1}^{2n-1}\omega_{i}\right)-{\mathrm{sgn}}\left(\sum_{i=1}^{2n-1}\omega_{i}\right)\right)$
$\sim$ $|\Omega|^{2n-1},$
which agrees with the scaling form of eqn. (37) when $m=1$.
The fermionic self-energy from $g$-coupling is given by
$\Sigma_{g}({\mathrm{i}}\omega)$ $=$ $g^{2}n\mathcal{N}^{2n-1}\int\left(\prod_{i=1}^{2n-1}\frac{d\omega_{i}}{2\pi}d\xi_{{\bm{k}}_{i}}\frac{1}{{\mathrm{i}}{\mathrm{sgn}}(\omega_{i})\Gamma/2-\xi_{{\bm{k}}_{i}}}\right)\int_{0}^{\infty}\frac{d^{d}{\bm{q}}}{(2\pi)^{2}}\frac{1}{{\bm{q}}^{2}+c_{B}(\omega-\sum_{i}^{2n-1}\omega_{i})^{\eta}}$ (68)
$\sim$ $g^{2}\int\left(\prod_{i=1}^{2n-1}d\omega_{i}{\mathrm{sgn}}(\omega_{i})\right)\left(\omega-\sum_{i}^{2n-1}\omega_{i}\right)^{\frac{\eta(d-2)}{2}}$
$\sim$ $g^{2}\omega^{2n-1+\frac{\eta(d-2)}{d}},$
which reproduces the scaling behaviour in eqn. (38).
We therefore conclude that eqns. (37) and (38) correctly capture the scaling behaviour of the self-energies even in the case $m=1$.
Appendix B Self-energies in Two Dimensions
This section reproduces the logarithmic scaling for $d=2$ via explicit calculation. This behaviour is not captured by the general result (29) used in the main text.
We begin by evaluating the integral
$J_{n}(\omega)$ $\equiv$ $\underset{\sum_{i}^{n}x_{i}\leq\omega}{\int}\left(\prod_{i=1}^{n}dx_{i}{\mathrm{sgn}}(x_{i})\right)\ln(\Big{|}\omega-\sum_{i}^{n}x_{i}\Big{|})$ (69)
$\sim$ $\underset{\sum_{i}^{n}x_{i}\leq\omega}{\int_{0}^{\infty}}\prod_{i=1}^{n}dx_{i}\ln(\Big{|}\omega-\sum_{i}^{n}x_{i}\Big{|}).$
Define $X\equiv\sum_{i}^{n}x_{i}$, and insert it with a delta function,
$J_{n}(\omega)$ $\sim$ $\underset{\sum_{i}^{n}x_{i}\leq\omega}{\int_{0}^{\infty}}\prod_{i=1}^{n}dx_{i}\ln(\Big{|}\omega-\sum_{i}^{n}x_{i}\Big{|})\int dX\delta(X-\sum_{i=1}^{n}x_{i})$ (70)
$=$ $\underset{0\leq X\leq\omega}{\int}dX\underset{\sum_{i=1}^{n-1}x_{i}\leq X}{\int}\prod_{i=1}^{n-1}dx_{i}\ln(\Big{|}\omega-X\Big{|})$
$\sim$ $\int_{0}^{\omega}dXX^{n-1}\ln(\omega-X).$
The dominant contribution comes from the region near $\omega$, where $\omega\ll 1$, so we expand the integrand as
$\ln(\omega-X)$ $=$ $\ln(\omega)+\ln(1-\frac{X}{\omega})\simeq\ln(\omega)-\frac{X}{\omega}+\mathcal{O}\left(\frac{X^{2}}{\omega^{2}}\right).$ (71)
Thus, we find the leading scaling behaviour
$J_{n}(\omega)$ $\sim$ $C_{1}\omega^{n}\ln(\omega)+C_{2}\omega^{n},$ (72)
where $C_{1}$ and $C_{2}$ are constants.
Now consider the full integral (35) in two dimensions:
$\mathcal{I}^{d=2}_{\alpha,\beta}({\mathrm{i}}x)$ $\equiv$ $\int_{-\infty}^{\infty}\left(\prod_{i=1}^{\alpha}\frac{d\omega_{i}}{2\pi}d\xi_{{\bm{k}}_{i}}\frac{1}{{\mathrm{i}}{\mathrm{sgn}}(\omega_{i})\Gamma/2-\xi_{{\bm{k}}_{i}}}\right)$ (73)
$\times\left(\prod_{j=1}^{\beta-1}\int_{-\infty}^{\infty}\frac{d\Omega_{j}}{2\pi}\int_{0}^{\infty}\frac{d^{2}{\bm{q}}_{j}}{(2\pi)^{2}}\frac{1}{|\Omega_{j}|^{\eta}+{\bm{q}}_{j}^{2}}\right)\int_{0}^{\infty}\frac{d^{2}{\bm{q}}_{\beta}}{(2\pi)^{2}}\frac{1}{|x+\sum_{j=1}^{\beta-1}\Omega_{j}+\sum_{i=1}^{2n}\omega_{i}|^{\eta}+{\bm{q}}_{\beta}^{2}}$
$\sim$ $(-{\mathrm{i}})^{\alpha}\int_{-\infty}^{\infty}\left(\prod_{i=1}^{\alpha}d\omega_{i}{\mathrm{sgn}}(\omega_{i})\right)\left(\prod_{j=1}^{\beta-1}\int_{-\infty}^{\infty}d\Omega_{j}\ln(\frac{\Lambda_{q}^{2}}{|\Omega_{j}|^{\eta}})\right)\ln(\frac{\Lambda_{q}^{2}}{|x+\sum_{j=1}^{\beta-1}\Omega_{j}+\sum_{i=1}^{2n}\omega_{i}|^{\eta}})$
$\sim$ $(-{\mathrm{i}})^{\alpha}\left(\prod_{j=1}^{\beta-1}\int_{-\infty}^{\infty}d\Omega_{j}\ln(\frac{\Lambda_{q}^{2}}{|\Omega_{j}|^{\eta}})\right)\left(\left(x+\sum_{j=1}^{\beta-1}\Omega_{j}\right)^{\alpha}\left(C_{1}\ln\left|x+\sum_{j=1}^{\beta-1}\Omega_{j}\right|+C_{2}\right)\right).$
Changing the variables $\{\Omega_{i}\}$ into $\{u_{i}\}$, such that
$\Omega_{j}\equiv x\cdot u_{j}\Leftrightarrow d\Omega_{j}=xdu_{j},$ (74)
one finds
$\mathcal{I}_{\alpha,\beta}^{d=2}({\mathrm{i}}x)$ $\sim$ $(-{\mathrm{i}})^{\alpha}x^{\alpha+\beta-1}\int\prod_{j=1}^{\beta-1}du_{j}\ln(x\cdot u_{j})\left(1+\sum_{j=1}^{\beta-1}u_{j}\right)^{\alpha}\ln(x\left(1+\sum_{j=1}^{\beta-1}u_{j}\right))$ (75)
$\sim$ $(-{\mathrm{i}})^{\alpha}x^{\alpha+\beta-1}\sum_{\iota=0}^{\beta}c_{\iota}(\ln x_{i})^{\iota}.$
The coefficient $c_{i}$ is fixed by explicit computation, but being $x$-independent, can remain unspecified in a qualitative analysis.
In particular, for the bosonic self-energy, $\alpha=2n$ and $\beta=m-1$, and for electronic self-energy, $\alpha=2n-1$ and $\beta=m$. We thus arrive at the general expressions in $d=2$,
$\Pi({\mathrm{i}}\Omega)-\Pi(0)$ $=$ $-g^{2}\Omega^{2n+m-2}\sum_{i=0}^{m-1}c_{i}(\ln(\Omega))^{i},$ (76)
$\Sigma({\mathrm{i}}\omega)$ $=$ $-{\mathrm{i}}g^{2}\omega^{2n+m-2}\sum_{i=0}^{m}c_{i}(\ln(\omega))^{i}.$ (77)
These results demonstrate explicitly that the logarithmic scaling, absent in the hypergeometric estimate, is correctly reproduced by direct evaluation in two dimensions.